OpenStudy (anonymous):

HELP Needed With # 2 & 4 #2 Take the log of both sides (x-1)log5 = x log 3 (x-1)0.698970004 = 0.477121255x ... now solve for x #4 check your arithmetic after you distributed (x-4)1.204119983=(3-x).477121254 1.204119983x - 4.816479931 = 1.431363764 -.4771212547x x = ? HELP ME PLEASE !

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=IIDTlgvo0DM

OpenStudy (anonymous):

#2: Ok, so you had the equation: \[5^{x-1}=3^x\] What they were asking and partially did is to solve for x So you take the log of both sides: \[log_{10}(5^{x-1})=log_{10}(3^x)\] And since \(\{log_a(b^c)=c\phantom{.}log_a(b)\}\), you obtain the following equation: \[(x-1)log_{10}(5)=xlog_{10}(3)\] Now you can distribute the left side to look like: \[xlog_{10}(5)-log_{10}(5)=xlog_{10}(3)\] And you gather "like" terms: \[xlog_{10}(5)-xlog_{10}(3)=log_{10}(5)\] Factor out an x: \[x\big(log_{10}(5)-log_{10}(3)\big)=log_{10}(5)\] And simplify since we know that \(\{log_c(a)-log_c(b)=log_c\left(\frac{a}{b}\right)\}\) \[x\Bigg(log_{10}\left(\frac{5}{3}\right)\Bigg)=log_{10}(5)\] And isolate for \(x\): \[x=\frac{log_{10}(5)}{log_{10}\left(\frac{5}{3}\right)}\dot{=}\phantom{0}3.15\]